# Multi-dimensional chaos I: Classical and quantum mechanics

### v1

@@ 1 Introduction
Chaotic processes often admit a description in terms of erratic functions of a certain continuous variable. Examples of such a behavior are the scattering angle as a function of the incident angle in a pinball experiment [1], the leaky torus phase shift [2] as a function of the wave-number and the decay amplitude of highly excited string state (HES) into two low-mass states or a scattering amplitude of a HES with three low-mass states as a function of an angle [3, 4].
In [5] and [6] we proposed a novel measure of chaotic behavior based on a map between the set of maxima of erratic scattering amplitudes and the eigenvalues of random matrices. The spacings between adjacent peaks and their ratios were shown to admit a gaussian $\beta$-ensemble distribution in a similar manner to the Wigner-Dyson map of the energy eigenvalues of chaotic complex nuclei and molecules [7].
@@
In the present work we analyze two-dimensional chaotic processes in the context of the two following frameworks: the classical [8] and the quantum [1] pinball scattering, and a toy model that generalizes the leaky torus scattering [2]. The latter takes the form of an electric potential produced by a set of charges located on a plane in three spatial dimensions.
For the classical pinball scattering we determine the scattering angle and number of collisions as a function of the incident angle and the impact parameter which is expressed also in terms of an angle. In fact the emerging patterns were studied intensively in the past, see [9] and references therein. To understand the patterns associated with chaotic processes, we compare the patterns of the three-disk chaotic cases to those of the non-chaotic scattering from two disks. In this way we identify the regions in the pattern that can be associated with the chaotic behavior.
It is a long-standing conjecture [10] that the eigenvalues of the $S$-matrices of quantum chaotic processes could be faithfully mapped into those of random matrices, specfically those of the circular ensembles of random unitary matrices  [^1] Chaotic $S$-matrix theory has been intensively investigated, see for instance [11, 12] and references therein. Surprisingly, we found out that for the fully symmetric setup, in spite of the fact that classical scattering is chaotic, the eigenvalues of the quantum $S$-matrix follow a non-chaotic distribution, namely a Poisson one. On the other hand for asymmetric setups and for a large enough wave-number $k$ the distribution of the $S$-matrix eigenvalues is the expected COE (circular orthogonal ensemble) distribution, associated with chaotic systems.
@@
Some similar measures in two dimensions has been analyzed in [13] and related works, in the context of dissipative quantum chaotic systems. There one encounters eigenvalues of the Hamiltonian which lie on the complex plane and can compare with Laguerre ensembles of random complex matrices. Since these methods do not easily generalize to higher dimensions, we will not utilize them in this work.
A commonly used quantity to identify chaos is the spectral form factor (SFF), see for instance [14]. In [15] an analogous quantity, dubbed the scattering form factor (ScFF), was introduced. It is straightforward to generalize this quantity for the vectors $\vec{\lambda}_{i},\ i=1,...N$ and correspondingly a vector of “times” $\vec{t}$.
@@
A cornerstone of quantum chaotic behavior is the ability to map certain discrete set of observables associated to a given system to the set of eigenvalues of random matrices [16]. This was done originally by Wigner and Dyson [17] and recently with the novel measure that we have proposed [5]. The Wigner-Dyson distribution of the (normalized) level spacings in the one-dimensional case is, explicitly:
@@
The distribution of consecutive spacing ratios, $r_{n}\equiv\delta_{n+1}/\delta_{n}$, is [18]:
@@
Section 4 is devoted to the analysis of the spacings between peaks of the amplitudes that depend on two variables. In 4.1 we first briefly review the landmark case of the phase shifts for the leaky torus [2] . We then propose a toy model that generalizes it with two random matrices. We consider a function of two random variables, which is the analog of the time delay of the leaky torus. This function can be also interpreted as the electric potential produced by a set of charges randomly located on a plane in three space dimensions. Section 4.2 is devoted to the analysis of two dimensional eigenvalues and their spacings. We propose four different measures for spacings in two (or higher) dimensions, including a definition to form a consecutive level spacings. These methods are applied to the toy model mentioned above in 4.3. We compute the corresponding distributions in two dimensions. We explain the resulting distribution by identifying the “effective repulsion” in the toy model. This is done by performing analytic calculations. Lastly in we consider a case where there is no disorder in the eigenvalues, in which they are taken to be integers but the pairing between them is performed with a permutation matrix. In this case also a GOE distribution is found for the distribution of the spacings. We end the analysis of the toy model by considering the two dimensional scattering form factor in 4.4.
@@ 2.2 Classical chaotic scattering
The classical chaotic behavior of the pinball scattering has been investigated thoroughly [8]. Several observables, such as the Lyapunov exponent(s), the Kolmogorov-Sinai entropy and other ones have been used as a measure of chaos in this context. In the present analysis we focus on two observables. One is the scattering angle, defined as the difference between the incoming and outgoing angles of the velocity. The second is the number of collisions $n_{c}$, closely related to the time the particle spends inside the system before escaping. We are interested in the dependence of these two observables on the two angles $\theta$ and $\phi$ defined above.
@@
A crucial feature of these functions is that they display a fractal, self-similar nature. Namely, when one zooms in on one of the chaotic regions, one always sees the same pattern of regions where the function is relatively smooth and points where the number of collisions is very large and the function is erratic. This pattern repeats indefinitely [8]. See figure 3. In the figure we zoom in on an erratic region in the plot of the scattering angle as a function of $\theta$, seeing the same structure repeating at all scales. One can see the same behavior in the scattering angle as a function of $\phi$.
@@ 2.3 Two-dimensional descriptions
The two dimensional patterns and their self-similar structure have been intensively studied in the past, see for instance [9] and references therein. In these references the scattering patterns were analyzed using the notion of Wada basins.
@@ 3.1 The exact quantization of the three-disk pinball system
The quantum version of the three-disk pinball scattering system was solved exactly by Gaspard and Rice [1]. The problem is simple to define. One has to solve the free wave-equation in the plane,
@@
In [1] the authors focus on the most symmetric case of three disks of equal radii centered on the corners of an equilateral triangle, the same as the classical setup we explored above. But this is not necessary, as the system can be solved for any general set-up, including $n$ disks of different radii and generic positions. We follow here the detailed review of [19], which contains the general solution. The analogous system of scattering from $n$ spheres in three dimensions has also been solved exactly using the same techniques [20].
@@
The $S$-matrix can be computed exactly using Green’s theorem to solve the equations for the wave-function with the correct boundary conditions. Here, we will not repeat the derivation and only write the final results. The details can be found in [1, 19].
@@
In principle, the angular momenta can take any integer value and the matrices are infinite dimensional. On the other hand, one can see that for a given configuration at finite $k$, there is a natural cutoff, that one can deduce from the behavior of the Bessel functions, when expanded for large order. For positive large $l$ and $x$ fixed [21],
@@
Since the arguments of the Bessel functions in our case are always $k$ times one of the characteristic lengths $L_{char}$ in the system ($R_{a}$, $L_{ab}$) this means that at fixed $k$ there is an effective finite dimension of the matrix $N\sim{\cal O}(kL_{char})$. In [1] it is pointed out that the only condition for the combination of Bessel functions appearing in $M_{am}{}^{a^{\prime}m^{\prime}}$ to be small at large enough $m$ or $m^{\prime}$ is that the disks do not overlap. This is always understood to be true in the pinball system.
@@ 3.2 The $S$-matrix of the pinball system
and the corresponding distribution of the normalized spacing ratios, $\tilde{r}_{n}=\min\left(\frac{\delta_{n+1}}{\delta_{n}},\frac{\delta_{n}}{\delta_{n+1}}\right)$ [18],
@@ 3.3 The quantum scattering amplitude
In [5, 6] we showed that for string scattering amplitudes involving highly-excited string states, one can find RMT distributions when looking at the spacings of consecutive peaks in the angular dependence of the amplitude. We would like to see if similar distributions emerge when looking at the angular dependence of the scattering amplitude in the pinball system.
We will consider the differential cross section, which is a function of the two angles, incoming and outgoing, as well as the energy, related to $k$. At fixed $k$, the differential cross-section given by $|f(\phi,\phi^{\prime})|^{2}$ is a complicated fluctuating function of the two angles, which we will examine in detail in the remainder of this section. Even though in this system we have already observed that there is a correspondence of the $S$-matrix to COE, we would like to examine if one could see this from examining directly the scattering amplitude, as we did for string amplitudes in [5, 6].
@@
In the classical case, there was no non-trivial dependence on the velocity, but in the quantum system the dependence of $k$ is significant. If we plot the total cross-section as a function of $k$, we can see a structure of resonances. However, the positions of peaks in the plot depend on whether one looks at a specific incoming angle or the average total cross-section, as can be seen in figure 12. The actual positions of the resonances (which have also an imaginary part) is most accurately determined from the locations of the singularities of the matrix $M$ (eq. ), and these also are quite regularly spaced, as noted already in [1]. Some long-lived resonances, i.e. with small imaginary part, can be observed as a small kink in the plot of the total cross section, for example around $k=12$ for the symmetric system plotted in figure 12.
@@ 3.4 Two dimensional description of the quantum differential cross section
  • There is no self-similarity. This is not surprising since there are other quantum chaotic systems that are not self similar, for instance the decays and scattering of highly excited string states [6, 22].
@@ 4 Modeling the spacings between peaks of scattering amplitudes in two dimensions
For the pinball system both the classical plots of the scattering angles and the quantum ones of the differential cross section as a function of the two angles admit a rich structure that include isolated peaks, minimum points, saddle points, ridges, valleys, etc. In our previous work [5, 6, 15], we emphasized the important role that the spacings between adjacent peaks of the scattering amplitudes play in chaotic scattering. We would like now to analyze this topic in the context of two-dimensional scattering. We then introduce and study models based on random matrices.
For this purpose we elaborate on toy models for scattering amplitudes with various types of ensembles of peaks. The models are based on a generalization of the phase shift of the so-called leaky torus. We start with a very brief description of the latter. We then introduce the two-dimensional toy models. Next we present the analysis of the two-dimensional spacings. We rely on random matrices for the locations of the peaks and determine various ways to analyze such systems and we end with the computation of the two-dimensional generalization of the scattering form factor (ScFF) introduced in [15].
@@ 4.1 The phase shifts of the leaky torus and the inspired toy model
A landmark model that displays chaotic behavior is scattering on the leaky torus. Originally proposed by Gutzwiller [2], the leaky torus geometry is constructed by taking the two-dimensional hyperbolic plane with the metric
@@
The Wigner time delay function is given in general by the determinant of the logarithmic derivative of the $S$-matrix, which here reduces to [^5] We mostly follow chapter 8 of [23].
@@ 6 A conjecture of a map to random tensor theory
An important measure of quantum chaos is the Wigner-Dyson [7] map of the energy eigenvalues to those of Gaussian random matrices [16]. In [24] we proposed a similar map replacing the energy eigenvalues with the peaks of scattering amplitudes. The latter took the form of erratic functions of a single variable. In the present investigation we have enlarged the scope and studied erratic functions of several variable. This raises a natural question whether one has to generalize the mathematical random object on the other side of the map, namely the random matrices. We focused here on the peaks of the multi-dimensional erratic functions. The latter are part of a richer set of ”topographic structures”. In section 4.2.1 it was shown that in the two dimensional case one can fix one variable and be left with the one dimensional case which can be mapped into a random matrix. If one asks what is a mathematical object that can be “fixed” and yield a random matrix, a natural answer is a 3-index (rank 3) random tensor, or random three-tensor for short.
@@ 6.1 From random tensors to distributions of spacings of peaks
The distributions of the eigenvectors were studied recently (see [25] and references therein). In particular real eigenvector/value distributions of Gaussian random tensors have been explicitly computed by expressing them as partition functions of quantum field theories with quartic interaction.
@@ 7 Summary
One of our aims was to introduce a measure to observe chaotic behavior directly from the scattering amplitudes, that would be applicable in cases where the full $S$-matrix is not known. The measure of chaotic behavior using the extrema of the scattering amplitude, introduced in [5, 6], was generalized to two and higher dimensions in this work.
To that end, we set out to analyze the spacings between peaks of amplitudes that depend on two variables. We first briefly reviewed the landmark case of the phase-shifts of the leaky torus [2], where the time delay function is a series of resonances located at the non-trivial zeros of the Riemann zeta function, and whose spacings are distributed as in the GUE.
@@
We defined and computed also the two-dimensional Scattering Form Factor (ScFF), introduced in [15] for scattering processes in analogy with the Spectral Form Factor (SFF). Unlike the simple SFF, the ScFF can be a function of multiple variables, associated with each of the kinematic variables in the problem. We have shown how the ScFF can shed light on the underlying distribution of the positions of extrema.
@@
  • In recent years increasing attention has been payed to the study of quantum chaos in QFTs in 1+1 dimension [26, 27, 28, 29, 30]. A natural question is to look for multi-dimensional chaotic processes in QFTs in 1+1 and higher dimensions.
  • An important tool that has emerged in the recent study of quantum chaos in quantum mechanical and QFT systems is Krylov complexity (see [31] and references therein). Can one develop a Lanczos method also for multi-dimensional chaos? In particular in [32] Krylov complexity methods were implemented for the measure of spacings between peaks of scattering amplitudes proposed in [5]. This probably can also be generalized to higher-dimensional chaotic behaviors.
  • Recently, in [33] and [34], the measure of [5] was used for computing correlators in quenched QFTs in particular in AdS. Again a generalization to higher dimensions and to other types of correlators should be possible.
  • We have started our journey of chaotic processes in string theory with the proposal for a quantitative measure [24] for the observed erratic behavior of scattering and decay amplitudes of HES states [3]. The latter are in fact generically functions of several kinematical parameters, and can be computed analytically at tree level. In a sequel to the present investigation we intend to present the results of the study of multi-dimensional chaotic behavior in such processes.

### v2

@@ 1 Introduction
Chaotic processes often admit a description in terms of erratic functions of a certain continuous variable. Examples of such a behavior are the scattering angle as a function of the incident angle in a pinball experiment [Gaspard:1989exq], the leaky torus phase shift [Gutzwiller:1983] as a function of the wave-number and the decay amplitude of highly excited string state (HES) into two low-mass states or a scattering amplitude of a HES with three low-mass states as a function of an angle [Gross:2021gsj, Rosenhaus:2021xhm].
In [Bianchi:2022mhs] and [Bianchi:2023uby] we proposed a novel measure of chaotic behavior based on a map between the set of maxima of erratic scattering amplitudes and the eigenvalues of random matrices. The spacings between adjacent peaks and their ratios were shown to admit a gaussian $\beta$-ensemble distribution in a similar manner to the Wigner-Dyson map of the energy eigenvalues of chaotic complex nuclei and molecules [Dyson:1962es].
@@
In the present work we analyze two-dimensional chaotic processes in the context of the two following frameworks: the classical [Gaspard:1989cla] and the quantum [Gaspard:1989exq] pinball scattering, and a toy model that generalizes the leaky torus scattering [Gutzwiller:1983]. The latter takes the form of an electric potential produced by a set of charges located on a plane in three spatial dimensions.
For the classical pinball scattering we determine the scattering angle and number of collisions as a function of the incident angle and the impact parameter which is expressed also in terms of an angle. In fact the emerging patterns were studied intensively in the past, see [Sweet:1999] and references therein. To understand the patterns associated with chaotic processes, we compare the patterns of the three-disk chaotic cases to those of the non-chaotic scattering from two disks. In this way we identify the regions in the pattern that can be associated with the chaotic behavior.
It is a long-standing conjecture [Blumel:1990zz] that the eigenvalues of the $S$-matrices of quantum chaotic processes could be faithfully mapped into those of random matrices, specifically those of the circular ensembles of random unitary matrices  [^1] Chaotic $S$-matrix theory has been intensively investigated, see for instance [Beenakker:1996esw, Haake:2010fgh] and references therein. Surprisingly, we found out that for the fully symmetric setup, in spite of the fact that classical scattering is chaotic, the eigenvalues of the quantum $S$-matrix follow a non-chaotic distribution, namely a Poisson one. On the other hand for asymmetric setups and for a large enough wave-number $k$ the distribution of the $S$-matrix eigenvalues is the expected COE (circular orthogonal ensemble) distribution, associated with chaotic systems. [^2] A similar observation was made in [Balasubramanian:2024ghv] regarding the level spacing statistics in triangular billiards systems.
@@
Some similar measures in two dimensions has been analyzed in [Sa:2020fpf] and related works, in the context of dissipative quantum chaotic systems. There one encounters eigenvalues of the Hamiltonian which lie on the complex plane and can compare with Laguerre ensembles of random complex matrices. Since these methods do not always generalize to higher dimensions, we will not utilize them directly in this work.
A commonly used quantity to identify chaos is the spectral form factor (SFF), see for instance [Cotler:2016fpe]. In [Bianchi:2024fsi] an analogous quantity, dubbed the scattering form factor (ScFF), was introduced. It is straightforward to generalize this quantity for the vectors $\vec{\lambda}_{i},\ i=1,...N$ and correspondingly a vector of “times” $\vec{t}$.
@@
A cornerstone of quantum chaotic behavior is the ability to map certain discrete set of observables associated to a given system to the set of eigenvalues of random matrices [Bohigas:1983er]. This was done originally by Wigner and Dyson [Dyson:1962b] and recently with the novel measure that we have proposed [Bianchi:2022mhs]. The Wigner-Dyson distribution of the (normalized) level spacings in the one-dimensional case is, explicitly:
@@
The distribution of consecutive spacing ratios, $r_{n}\equiv\delta_{n+1}/\delta_{n}$, is [Atas:2013dis]:
@@
Section 4 is devoted to the analysis of the spacings between peaks of the amplitudes that depend on two variables. In 4.1 we first briefly review the landmark case of the phase shifts for the leaky torus [Gutzwiller:1983] . We then propose a toy model that generalizes it with two random matrices. We consider a function of two random variables, which is the analog of the time delay of the leaky torus. This function can be also interpreted as the electric potential produced by a set of charges randomly located on a plane in three space dimensions. Section 4.2 is devoted to the analysis of two dimensional eigenvalues and their spacings. We propose four different measures for spacings in two (or higher) dimensions, including a definition to form a consecutive level spacings. These methods are applied to the toy model mentioned above in 4.3. We compute the corresponding distributions in two dimensions. We explain the resulting distribution by identifying the “effective repulsion” in the toy model. This is done by performing analytic calculations. Lastly in we consider a case where there is no disorder in the eigenvalues, in which they are taken to be integers but the pairing between them is performed with a permutation matrix. In this case also a GOE distribution is found for the distribution of the spacings. We end the analysis of the toy model by considering the two dimensional scattering form factor in 4.4.
@@ 2.2 Classical chaotic scattering
The classical chaotic behavior of the pinball scattering has been investigated thoroughly [Gaspard:1989cla]. Several observables, such as the Lyapunov exponent(s), the Kolmogorov-Sinai entropy and other ones have been used as a measure of chaos in this context. In the present analysis we focus on two observables. One is the scattering angle, defined as the difference between the incoming and outgoing angles of the velocity. The second is the number of collisions $n_{c}$, closely related to the time the particle spends inside the system before escaping. We are interested in the dependence of these two observables on the two angles $\theta$ and $\phi$ defined above.
@@
A crucial feature of these functions is that they display a fractal, self-similar nature. Namely, when one zooms in on one of the chaotic regions, one always sees the same pattern of regions where the function is relatively smooth and points where the number of collisions is very large and the function is erratic. This pattern repeats indefinitely [Gaspard:1989cla]. See figure 3. In the figure we zoom in on an erratic region in the plot of the scattering angle as a function of $\theta$, seeing the same structure repeating at all scales. One can see the same behavior in the scattering angle as a function of $\phi$.
@@ 2.3 Two-dimensional descriptions
The two dimensional patterns and their self-similar structure have been intensively studied in the past, see for instance [Sweet:1999] and references therein. In these references the scattering patterns were analyzed using the notion of Wada basins.
@@ 3.1 The exact quantization of the three-disk pinball system
The quantum version of the three-disk pinball scattering system was solved exactly by Gaspard and Rice [Gaspard:1989exq]. The problem is simple to define. One has to solve the free wave-equation in the plane,
@@
In [Gaspard:1989exq] the authors focus on the most symmetric case of three disks of equal radii centered on the corners of an equilateral triangle, the same as the classical setup we explored above. But this is not necessary, as the system can be solved for any general set-up, including $n$ disks of different radii and generic positions. We follow here the detailed review of [Wirzba:1997hy], which contains the general solution. The analogous system of scattering from $n$ spheres in three dimensions has also been solved exactly using the same techniques [Henseler:1997].
@@
The $S$-matrix can be computed exactly using Green’s theorem to solve the equations for the wave-function with the correct boundary conditions. Here, we will not repeat the derivation and only write the final results. The details can be found in [Gaspard:1989exq, Wirzba:1997hy].
@@
In principle, the angular momenta can take any integer value and the matrices are infinite dimensional. On the other hand, one can see that for a given configuration at finite $k$, there is a natural cutoff, that one can deduce from the behavior of the Bessel functions, when expanded for large order. For positive large $l$ and $x$ fixed [NIST:DLMF],
@@
Since the arguments of the Bessel functions in our case are always $k$ times one of the characteristic lengths $L_{char}$ in the system ($R_{a}$, $L_{ab}$) this means that at fixed $k$ there is an effective finite dimension of the matrix $N\sim{\cal O}(kL_{char})$. In [Gaspard:1989exq] it is pointed out that the only condition for the combination of Bessel functions appearing in $M_{am}{}^{a^{\prime}m^{\prime}}$ to be small at large enough $m$ or $m^{\prime}$ is that the disks do not overlap. This is always understood to be true in the pinball system.
@@ 3.2 The $S$-matrix of the pinball system
and the corresponding distribution of the normalized spacing ratios, $\tilde{r}_{n}=\min\left(\frac{\delta_{n+1}}{\delta_{n}},\frac{\delta_{n}}{\delta_{n+1}}\right)$ [Atas:2013dis],
@@ 3.3 The quantum scattering amplitude
In [Bianchi:2022mhs, Bianchi:2023uby] we showed that for string scattering amplitudes involving highly-excited string states, one can find RMT distributions when looking at the spacings of consecutive peaks in the angular dependence of the amplitude. We would like to see if similar distributions emerge when looking at the angular dependence of the scattering amplitude in the pinball system.
We will consider the differential cross section, which is a function of the two angles, incoming and outgoing, as well as the energy, related to $k$. At fixed $k$, the differential cross-section given by $|f(\phi,\phi^{\prime})|^{2}$ is a complicated fluctuating function of the two angles, which we will examine in detail in the remainder of this section. Even though in this system we have already observed that there is a correspondence of the $S$-matrix to COE, we would like to examine if one could see this from examining directly the scattering amplitude, as we did for string amplitudes in [Bianchi:2022mhs, Bianchi:2023uby].
@@
In the classical case, there was no non-trivial dependence on the velocity, but in the quantum system the dependence of $k$ is significant. If we plot the total cross-section as a function of $k$, we can see a structure of resonances. However, the positions of peaks in the plot depend on whether one looks at a specific incoming angle or the average total cross-section, as can be seen in figure 12. The actual positions of the resonances (which have also an imaginary part) is most accurately determined from the locations of the singularities of the matrix $M$ (eq. ), and these also are quite regularly spaced, as noted already in [Gaspard:1989exq]. Some long-lived resonances, i.e. with small imaginary part, can be observed as a small kink in the plot of the total cross section, for example around $k=12$ for the symmetric system plotted in figure 12.
@@ 3.4 Two dimensional description of the quantum differential cross section
  • There is no self-similarity. This is not surprising since there are other quantum chaotic systems that are not self similar, for instance the decays and scattering of highly excited string states [Bianchi:2023uby, Hashimoto:2022bll].
@@ 4 Modeling the spacings between peaks of scattering amplitudes in two dimensions
For the pinball system both the classical plots of the scattering angles and the quantum ones of the differential cross section as a function of the two angles admit a rich structure that include isolated peaks, minimum points, saddle points, ridges, valleys, etc. In our previous work [Bianchi:2022mhs, Bianchi:2023uby, Bianchi:2024fsi], we emphasized the important role that the spacings between adjacent peaks of the scattering amplitudes play in chaotic scattering. We would like now to analyze this topic in the context of two-dimensional scattering. We then introduce and study models based on random matrices.
For this purpose we elaborate on toy models for scattering amplitudes with various types of ensembles of peaks. The models are based on a generalization of the phase shift of the so-called leaky torus. We start with a very brief description of the latter. We then introduce the two-dimensional toy models. Next we present the analysis of the two-dimensional spacings. We rely on random matrices for the locations of the peaks and determine various ways to analyze such systems and we end with the computation of the two-dimensional generalization of the scattering form factor (ScFF) introduced in [Bianchi:2024fsi].
@@ 4.1 The phase shifts of the leaky torus and the inspired toy model
A landmark model that displays chaotic behavior is scattering on the leaky torus. Originally proposed by Gutzwiller [Gutzwiller:1983], the leaky torus geometry is constructed by taking the two-dimensional hyperbolic plane with the metric
@@
The Wigner time delay function is given in general by the determinant of the logarithmic derivative of the $S$-matrix, which here reduces to [^6] We mostly follow chapter 8 of [Hurt:1997quantum].
@@ 6 A conjecture of a map to random tensor theory
An important measure of quantum chaos is the Wigner-Dyson [Dyson:1962es] map of the energy eigenvalues to those of Gaussian random matrices [Bohigas:1983er]. In [Bianchi:2022qph] we proposed a similar map replacing the energy eigenvalues with the peaks of scattering amplitudes. The latter took the form of erratic functions of a single variable. In the present investigation we have enlarged the scope and studied erratic functions of several variable. This raises a natural question whether one has to generalize the mathematical random object on the other side of the map, namely the random matrices. We focused here on the peaks of the multi-dimensional erratic functions. The latter are part of a richer set of ”topographic structures”. In section 4.2.1 it was shown that in the two dimensional case one can fix one variable and be left with the one dimensional case which can be mapped into a random matrix. If one asks what is a mathematical object that can be “fixed” and yield a random matrix, a natural answer is a 3-index (rank 3) random tensor, or random three-tensor for short.
@@ 6.1 From random tensors to distributions of spacings of peaks
The distributions of the eigenvectors were studied recently (see [Sasakura:2023crd] and references therein). In particular real eigenvector/value distributions of Gaussian random tensors have been explicitly computed by expressing them as partition functions of quantum field theories with quartic interaction.
@@ 7 Summary
One of our aims was to introduce a measure to observe chaotic behavior directly from the scattering amplitudes, that would be applicable in cases where the full $S$-matrix is not known. The measure of chaotic behavior using the extrema of the scattering amplitude, introduced in [Bianchi:2022mhs, Bianchi:2023uby], was generalized to two and higher dimensions in this work.
To that end, we set out to analyze the spacings between peaks of amplitudes that depend on two variables. We first briefly reviewed the landmark case of the phase-shifts of the leaky torus [Gutzwiller:1983], where the time delay function is a series of resonances located at the non-trivial zeros of the Riemann zeta function, and whose spacings are distributed as in the GUE.
@@
We defined and computed also the two-dimensional Scattering Form Factor (ScFF), introduced in [Bianchi:2024fsi] for scattering processes in analogy with the Spectral Form Factor (SFF). Unlike the simple SFF, the ScFF can be a function of multiple variables, associated with each of the kinematic variables in the problem. We have shown how the ScFF can shed light on the underlying distribution of the positions of extrema.
@@
  • In recent years increasing attention has been payed to the study of quantum chaos in QFTs in 1+1 dimension [Brandino:2010sv, Srdinsek:2020bpq, Delacretaz:2022ojg, Negro:2022hno, Sonnenschein:2025jlc]. A natural question is to look for multi-dimensional chaotic processes in QFTs in 1+1 and higher dimensions.
  • An important tool that has emerged in the recent study of quantum chaos in quantum mechanical and QFT systems is Krylov complexity (see [Nandy:2024evd, Rabinovici:2025otw] and references therein). Can one develop a Lanczos method also for multi-dimensional chaos? In particular in [Bhattacharya:2024szw] Krylov complexity methods were implemented for the measure of spacings between peaks of scattering amplitudes proposed in [Bianchi:2022mhs]. This probably can also be generalized to higher-dimensional chaotic behaviors.
  • Recently, in [Ageev:2025yiq] and [Ageev:2025iiy], the measure of [Bianchi:2022mhs] was used for computing correlators in quenched QFTs in particular in AdS. Again a generalization to higher dimensions and to other types of correlators should be possible.
  • We have started our journey of chaotic processes in string theory with the proposal for a quantitative measure [Bianchi:2022qph] for the observed erratic behavior of scattering and decay amplitudes of HES states [Gross:2021gsj]. The latter are in fact generically functions of several kinematical parameters, and can be computed analytically at tree level. In a sequel to the present investigation we intend to present the results of the study of multi-dimensional chaotic behavior in such processes.